OpenStudy (anonymous):

If a merchant has two types of tea, one work $2.70 per kilogram and the other worth $3.00 per kilogram, how many kilograms of each type should the merchant use in order to produce 30 kilograms of a blend that is worth $2.95 per kilogram?

OpenStudy (dumbcow):

A + B = 30 2.7A + 3B = 2.95*30 use substitution to find A and B

OpenStudy (anonymous):

dumb cow is not that dumb after all. huh

OpenStudy (dumbcow):

haha thanks dotdotdog can you figure it out now

OpenStudy (anonymous):

ya ur medal speaks 4 u...